OpenStudy (anonymous):

7x2 + 12x = 4 A. -2/7, -2 B. 2/7 C. 2/7, -2 D. -2 E. None of the above

OpenStudy (anonymous):

7x^2 +12x = 4 7x^2 +12x -4 =0 x^2 +12x/7 - 4/7 = 0 solve the quadratic

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Correct :)